OpenStudy (anonymous):

Classify the conic section and write its equation in standard form. 1) y2 - 8y - 8x = -24 2.) x2 + y2- 4x + 2y - 11 = 0 3). 4x2 + y2 - 8x - 8 = 0 4) 4x2 - 25y2 - 100y = 0

OpenStudy (amistre64):

youll most likely have to complete some squares

OpenStudy (amistre64):

16-16 .... (y-4)^2 = 8x-8 (y-4)^2 = 8(x-1) 1/8 (y-4)^2 + 1= x seems rather parabolic to me